OpenStudy (danielajohana):

Write the expression in factored form as an algebraic expression of a single trigonometric function. (sin^2 (x) -1)/ 1+sinx

OpenStudy (danielajohana):

It looks like this:\[\frac{ \sin ^{2}x -1 }{ 1+\sin x}\]